# On the First Passage Times of Branching Random Walks in $\mathbb{R}^{d}$

### v1

@@ 3.2 Finer asymptotics for non-spherically symmetric jumps
It holds that $\widehat{c}_{1}=c_{1}$ for all $\rho>1$ if and only if $\mathbb{E}[\xi_{i}\mid\xi_{1}]=0$ for all $2\leqslant i\leqslant d$, where $\xi_{i}$ denotes the $i$-th coordinate of the $\mathbb{R}^{d}$-valued random variable ${\bm{\xi}}$.
@@
$\sup_{\lambda_{1}\in\mathbb{R}}\left(\lambda_{1}c_{1}-\mathbb{E}[e^{\lambda_{1}\xi_{1}}]\right)=\sup_{\lambda_{1},\dots,\lambda_{d}\in\mathbb{R}}\left(\lambda_{1}c_{1}-\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]\right).$
@@
$\forall\lambda_{1}\in\mathbb{R},\leavevmode\nobreak\ \inf_{\lambda_{2},\dots,\lambda_{d}\in\mathbb{R}}\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]=\mathbb{E}[e^{\lambda_{1}\xi_{1}}].$
[…]
@@
    the law of ${\bm{\xi}}=(\xi_{1},\dots,\xi_{d})$ satisfies $\mathbb{E}[\xi_{i}\mid\xi_{1}]=0$ a.s. for all $2\leqslant j\leqslant d$, and satisfies (A2) and (A4);
@@ 4 Numerical analysis
In this section, we test Theorem 3 numerically. Since tracking the locations of all particles in a BRW is time-consuming and memory-intensive, we introduce a path-purging algorithm that removes particles that are highly unlikely to have descendants that realize the first passage time. [^12] An efficient algorithm that simulates the tip of one-dimensional BBM has been proposed in [18]. Unfortunately, the arguments therein do not seem to carry over to general BRWs in higher dimensions. A preliminary version of the algorithm for spherically symmetric jumps can be found in Section 2.3 of [54].
@@
Finally, we note that the main result of [30] shows that there rarely exists an infinite path in a one-dimensional BRW that always stays above the curve $n\mapsto(c_{1}-\varepsilon)n$ for $\varepsilon>0$ small. While this suggests that our path purging algorithm cannot be backed up by theory for a fixed $q_{c}$ as $x\to\infty$, the simulations work effectively well in the regime where $q_{c}$ is quite large compared to $x$.
@@ 4.1 Uniform distribution on the sphere
[…]
@@
Figure 5: The numerical 3D BRW obtained $c_{1}$ compared against the reference calculations at different $p_{3}$ for the (a) classical BRW and (b) delayed branching BRW with $p_{0}=0.004$. The mean of the FPT at each $x$ is computed from $2000$ samples.
@@
[…]
Figure 6 shows the FPT distribution at different $x$. Figure 6 shows the linear increase in the mean FPT (negligible logarithmic effect), whereas the Figure 6 validates the asymptotic in (55) by showing a negligible change in the standard deviation of FPT distribution with the increase in the offset distance, $x$.
@@ 4.2 Gaussian distributions
Suppose now that the jump ${\bm{\xi}}$ is centered Gaussian with a positive definite covariance $\Sigma$. We have $\Lambda(\bm{\lambda})=\bm{\lambda}^{\top}\Sigma\bm{\lambda}/2$ and $I(\mathbf{x})=\mathbf{x}^{\top}\Sigma^{-1}\mathbf{x}/2$. In particular, $\mathbf{c}_{2}=\nabla I(x/n,\mathbf{0})=((\Sigma^{-1})_{11}x/n,\dots,(\Sigma^{-1})_{1d}x/n)$ and $\Lambda(\mathbf{c}_{2})=(x/n)^{2}(\Sigma^{-1})_{11}/2$. By (60), the normal vector $\mathbf{n}$ is a constant multiple of $\Sigma^{-1}\mathbf{e}_{1}$, where $\mathbf{e}_{1}=(1,\mathbf{0})\in\mathbb{R}^{d}$. If $\Sigma$ is diagonal (independent jumps), the normal vector $\mathbf{n}$ is then parallel to the direction of the termination site $\mathbf{x}$. This is the case in our previous work [54] as opposed to the general case of non-spherically symmetric and dependent jumps that we will discuss in this section.
We look at three specific examples: (i) symmetric and independent (S.I.) jumps (presented earlier [54]), (ii) non-symmetric and independent (N-S.I.) jumps, and (iii) non-symmetric and dependent (N-S.D.) jumps. The corresponding covariances for the three cases are given by
@@
We follow the same procedures as Section 4.1 that estimate the statistic $\widehat{c}_{1}$ in (7) through fitting (61). Note that at any given $p_{3}$, the $\widehat{c}_{1}$ is a function of $(\Sigma^{-1})_{11}$. As a result, we expect the $c_{1}$ of the Gaussian BRW with S.I. and N-S.I. jumps to be identical as confirmed by our numerical results in Figures 7 and 7. We use $p_{0}=0$ for the comparison between the numerical and theoretical results in this section.
@@ 5 Concluding remarks
In this paper, we studied the first passage times of an $\mathbb{R}^{d}$-valued branching random walk (BRW) to a shifted unit ball. The first passage time asymptotics consists of a linear term and a logarithm correction term, as a function of the Euclidean distance of the target ball from the origin. We discussed extensions of this result to the delayed branching BRW model. As an immediate application, we obtained theoretical predictions of the shortest path statistics for polymer networks consisting of long chains and random cross-links. We conclude with a conjecture on finer asymptotics of the first passage times.
The first passage times of the branching Brownian motion, the continuous-time sibling of the BRW, possess better tractability due to its connection to the Fisher-KPP equation. It is proven in Section 4.1 of [54] the finer asymptotic for the first passage times of a standard branching Brownian motion that
@@
where the $O_{\mathbb{P}}(1)$ is tight. We refer the readers to [54] for the proof and a detailed discussion. In view also of the tightness result of Theorem 2, it is natural to postulate the following conjecture for BRW. Recall the constant $\widehat{c}_{1}$ from Section 1.2.
@@
acknowledgement
We thank Amir Dembo, Haotian Gu, and Alexandra Stavrianidi for helpful discussions. The material in this paper is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397. Additional support is gratefully acknowledged from NSF 1915967, 2118199, 2229012, 2312204.
@@ acknowledgement
Appendix A Large deviation estimates
@@
Consider $d\geqslant 2$. Suppose that the radial component $R$ of the spherically symmetric random variable ${\bm{\xi}}$ has exponential moments and $\mathbb{P}(R=0)<1$. Let $\bm{\lambda}\in\mathbb{R}^{d}$ be such that $\mathbb{E}_{\mathbb{P}}[e^{\bm{\lambda}\cdot\mathbf{X}}]$ exists and define a probability measure $\mathbb{Q}$ by $\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}(\mathbf{x}):=e^{\bm{\lambda}\cdot\mathbf{x}}/\mathbb{E}_{\mathbb{P}}[e^{\bm{\lambda}\cdot\mathbf{X}}]$ where $\mathbb{P}$ is the law of $\mathbf{X}$. Then the law of ${\bm{\xi}}$ under $\mathbb{Q}$ satisfies the strong non-lattice property
@@
Moreover, consider an i.i.d. sequence $\{\mathbf{X}_{n}\}_{n\geqslant 1}$ with the same distribution as ${\bm{\xi}}$ under $\mathbb{Q}$ with a positive definite covariance matrix $V$ and suppose that $\mathbb{E}_{\mathbb{Q}}[\left\lVert{\bm{\xi}}\right\rVert^{3}]<\infty$. Then
$\sup_{B}\bigg{|}\frac{1}{\sqrt{n}}\sum_{j=1}^{n}(\mathbf{X}_{j}-\mathbb{E}_{\mathbb{Q}}[\mathbf{X}_{j}])-\mu_{\mathbf{0},V}(B)\bigg{|}=o(1),$
where the supremum is taken over all boxes $B\subseteq\mathbb{R}^{d}$ and $\mu_{\mathbf{0},V}$ is the Gaussian measure on $\mathbb{R}^{d}$ with mean $\mathbf{0}$ and covariance $V$.
@@
By Theorem 2.10 of [26], for $y\in[-1,1]$ and some $C(d)\geqslant 0$,
@@
$z_{R_{1},R_{2}}(r,\bm{\theta})=C(d)r\left(1-\left(\frac{r^{2}-R_{1}^{2}-R_{2}^{2}}{2R_{1}R_{2}}\right)^{2}\right)^{\frac{d-3}{2}},\leavevmode\nobreak\ r\in[|R_{1}-R_{2}|,R_{1}+R_{2}],\leavevmode\nobreak\ \bm{\theta}\in\mathbb{S}^{d-1}.$
Therefore, the conditional law of ${\bm{\zeta}}$ under $\mathbb{P}$, and hence also the unconditional law under $\mathbb{Q}$, has a density in $\mathbb{R}^{d}$. If $\mathbb{P}(R=0)>0$, the laws have densities except for an atom at zero, and the densities are non-trivial since $\mathbb{P}(R=0)<1$. By Lemma 4 in Section XV.4 of [27], (63) holds with ${\bm{\zeta}}$ in place of ${\bm{\xi}}$, and hence (63) holds since ${\bm{\zeta}}$ is a sum of two i.i.d. copies of ${\bm{\xi}}$. The final claim follows from Theorem 2(b) of [9]. ∎
@@
Let $\Lambda(\lambda)=\log\phi_{\xi}(\lambda)=\log\mathbb{E}[e^{\lambda\xi}]$. It is clear from (A2) that $\xi$ is non-lattice. The Bahadur-Rao theorem (Theorem 3.7.4 of [22]) then implies that uniformly in $c\in[-C\log n,C\log n]$,
$\mathbb{P}(X\geqslant m_{n}+c)\asymp\frac{e^{-nI((m_{n}+c)/n)}}{\sqrt{n}}\asymp\frac{e^{-nI(c_{1})-(c-3\log n/2)I^{\prime}(c_{1})}}{\sqrt{n}}=n\rho^{-n}e^{-c_{2}c}.$ (64)
@@
satisfies that under $\mathbb{Q}$, $\{X_{i}\}_{i\in\mathbb{N}}$ are i.i.d. with mean $c_{1}$. The random variable $(\widetilde{X}_{i},\mathbf{Y}_{i})$ is then centered under $\mathbb{Q}$ where $\widetilde{X}_{i}=X_{i}-c_{1}$. Define $\widetilde{X}:=\widetilde{X}_{1}+\dots+\widetilde{X}_{n}$. It follows that
$\hskip 14.22636pt\mathbb{P}(X\geqslant m_{n}+c,\,\mathbf{Y}\in B_{\mathbf{0}}(u(n)))$
$=e^{n\Lambda(c_{2})}\mathbb{E}_{\mathbb{Q}}\left[e^{-c_{2}X}{\mathbbm{1}}_{\{X\geqslant m_{n}+c,\,\mathbf{Y}\in B_{\mathbf{0}}(u(n))\}}\right]$
$\asymp e^{n\Lambda(c_{2})-m_{n}c_{2}}\sum_{j=c+1}^{\infty}e^{-c_{2}j}\mathbb{Q}\left(\widetilde{X}\in\Big{[}j-1-\frac{3}{2c_{2}}\,\log n,j-\frac{3}{2c_{2}}\,\log n\Big{)},\,\mathbf{Y}\in B_{\mathbf{0}}(u(n))\right)$
$=\rho^{-n}n^{3/2}\sum_{j=c+1}^{\infty}e^{-c_{2}j}\mathbb{Q}\left(\frac{(\widetilde{X},\mathbf{Y})}{\sqrt{n}}\in\Big{[}\frac{j-1-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}},\frac{j-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}}\Big{)}\times B_{\mathbf{0}}\Big{(}\frac{u(n)}{\sqrt{n}}\Big{)}\right).$
@@
$\mathbb{Q}\left(\frac{(\widetilde{X},\mathbf{Y})}{\sqrt{n}}\in E_{j,n}\right)=(1+o(1))\mu_{\mathbf{0},V}\left(E_{j,n}\right)\asymp u(n)^{d-1}n^{-d/2},$ (66)
@@
$\mathbb{P}(\mathbf{Y}\in B_{\mathbf{0}}(u(n))\mid X\geqslant m_{n})$ $\ll(n\rho^{-n}e^{-c_{2}c})^{-1}\rho^{-n}n^{3/2}\sum_{j=c+1}^{\infty}e^{-c_{2}j}u(n)^{d-1}n^{-d/2}$
@@
$\hskip 14.22636pt\mathbb{P}(\mathbf{Y}\in B_{\mathbf{y}}(1),\leavevmode\nobreak\ X\in[m_{n}+c,m_{n}+c+a(n)])$
$\asymp\rho^{-n}n^{3/2}\sum_{j=c+1}^{c+a(n)}e^{-c_{2}j}\mathbb{Q}\left(\frac{(\widetilde{X},\mathbf{Y})}{\sqrt{n}}\in\Big{[}\frac{j-1-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}},\frac{j-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}}\Big{)}\times B_{\frac{\mathbf{y}}{\sqrt{n}}}\Big{(}\frac{1}{\sqrt{n}}\Big{)}\right)$
@@
This proves (21). The same computation also yields
$\mathbb{P}(\mathbf{Y}\in B_{\mathbf{0}}(K\sqrt{n}),\leavevmode\nobreak\ X\geqslant m_{n}+c)$ $\ll\rho^{-n}n^{3/2}n^{-1/2}=n\rho^{-n},$

### v2

@@ 3.2 Finer asymptotics for non-spherically symmetric jumps
It holds that $\widehat{c}_{1}=c_{1}$ for all $\rho>1$ if and only if $\mathbb{E}[{\xi^{(i)}}\mid{\xi^{(1)}}]=0$ for all $2\leqslant i\leqslant d$, where $\xi^{(i)}$} denotes the $i$-th coordinate of the $\mathbb{R}^{d}$-valued random variable $\bm{\xi}$}.
@@
$\sup_{\lambda_{1}\in\mathbb{R}}\left(\lambda_{1}c_{1}-\mathbb{E}[e^{\lambda_{1}{\xi^{(1)}}}]\right)=\sup_{\lambda_{1},\dots,\lambda_{d}\in\mathbb{R}}\left(\lambda_{1}c_{1}-\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]\right).$
@@
$\forall\lambda_{1}\in\mathbb{R},\penalty 10000\ \inf_{\lambda_{2},\dots,\lambda_{d}\in\mathbb{R}}\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]=\mathbb{E}[e^{\lambda_{1}{\xi^{(1)}}}].$
By taking partial derivatives with respect to $\lambda_{i}$ at $\lambda_{i}=0,\penalty 10000\ 2\leqslant i\leqslant d$, we obtain that for each $2\leqslant i\leqslant d,\penalty 10000\ \mathbb{E}[{\xi^{(i)}}e^{\lambda_{1}{\xi^{(1)}}}]=0$. Since this holds for all $\lambda_{1}\in\mathbb{R}$, by the uniqueness of the Laplace transform, we must have $\mathbb{E}[{\xi^{(i)}}\mid{\xi^{(1)}}]=0$. The reverse direction is obvious by Jensen’s inequality. ∎
@@
    the law of ${\bm{\xi}}=({\xi^{(1)}},\dots,{\xi^{(d)}})$ satisfies $\mathbb{E}[{\xi^{(i)}}\mid{\xi^{(1)}}]=0$ a.s. for all $2\leqslant j\leqslant d$, and satisfies (A2) and (A4);
@@ 4 Numerical analysis
In this section, we test Theorem 3 numerically. Since tracking the locations of all particles in a BRW is time-consuming and memory-intensive, we introduce a path-purging algorithm that removes particles that are highly unlikely to have descendants that realize the first passage time. [^13] An efficient algorithm that simulates the tip of one-dimensional BBM has been proposed in [brunet2020generate]. Unfortunately, the arguments therein do not seem to carry over to general BRWs in higher dimensions. A preliminary version of the algorithm for spherically symmetric jumps can be found in Section 2.3 of [zhang2024modeling].
@@
Finally, we note that the main result of [gantert2011asymptotics] shows that there rarely exists an infinite path in a one-dimensional BRW that always stays above the curve $n\mapsto(c_{1}-\varepsilon)n$ for $\varepsilon>0$ small. While this suggests that our path purging algorithm cannot be backed up by theory for a fixed $q_{c}$ as $x\to\infty$, the simulations work effectively well in the regime where $q_{c}$ is quite large compared to $x$.
@@ 4.1 Uniform distribution on the sphere
Consider the case when $\bm{\xi}$} is uniformly distributed in $\mathbb{S}^{2}$, i.e., we are in the spherically symmetric case. The statistic that we compare is the coefficient of the linear term in the estimation of $\tau_{x}$ in (4). For any given $p_{3}\in\{0.05\ell:1\leqslant\ell\leqslant 19\}$, in the numerical BRW simulation, we calculate the FPT distribution at different offset distances, $x\in\{0.25L_{x},\,0.5L_{x},\,0.75L_{x},\,L_{x}\}$, where $L_{x}=65.5$. The mean FPT $\mathbb{E}[\tau_{x}]$ is then fit to a function of the form
@@
Figure 5: The numerical 3D BRW obtained $c_{1}$ compared against the reference calculations across different $p_{3}$ values for the (a) classical BRW and (b) delayed branching BRW with $p_{0}=0.004$. The mean of the FPT at each $x$ is computed from $2000$ samples.
@@
Figure 6: (a) The FPT distributions at different values of $x$ for the classical BRW. The (b) mean and (c) standard deviation of the SP as a function of $x$, respectively. The data points correspond to $x=0.1L_{x},\,0.15L_{x},\,0.2L_{x},\,0.25L_{x},\,0.5L_{x},\,0.75L_{x},\,L_{x}$ and $p_{3}=0.0856$ for the classical BRW, where $L_{x}=65.5$. The mean of the FPT at each $x$ is computed from $500$ samples.
Figure 6 shows the FPT distribution at different $x$. Figure 6 shows the linear increase in the mean FPT (negligible logarithmic effect), whereas Figure 6 validates the asymptotic in (55) by showing a negligible change in the standard deviation of the FPT distribution with increasing offset distance, $x$.
@@ 4.2 Gaussian distributions
Suppose now that the jump $\bm{\xi}$} is centered Gaussian with a positive definite covariance $\Sigma$. We have $\Lambda(\bm{\lambda})=\bm{\lambda}^{\top}\Sigma\bm{\lambda}/2$ and $I(\mathbf{x})=\mathbf{x}^{\top}\Sigma^{-1}\mathbf{x}/2$. In particular, $\mathbf{c}_{2}=\nabla I(x/n,\mathbf{0})=((\Sigma^{-1})_{11}x/n,\dots,(\Sigma^{-1})_{1d}x/n)$ and $\Lambda(\mathbf{c}_{2})=(x/n)^{2}(\Sigma^{-1})_{11}/2$. By (60), the normal vector $\mathbf{n}$ is a constant multiple of $\Sigma^{-1}\mathbf{e}_{1}$, where $\mathbf{e}_{1}=(1,\mathbf{0})\in\mathbb{R}^{d}$. If $\Sigma$ is diagonal (independent jumps), the normal vector $\mathbf{n}$ is then parallel to the direction of the termination site $\mathbf{x}$. This is the case in our previous work [zhang2024modeling] as opposed to the general case of non-spherically symmetric and dependent jumps that we will discuss in this section.
We look at three specific examples: (i) symmetric and independent (S.I.) jumps (presented earlier [zhang2024modeling]), (ii) non-symmetric and independent (N-S.I.) jumps, and (iii) non-symmetric and dependent (N-S.D.) jumps. The corresponding covariances for the three cases are given by
@@
We follow the same procedures as Section 4.1 that estimate the statistic $\widehat{c}_{1}$ in (7) through fitting (61). Note that at any given $p_{3}$, the $\widehat{c}_{1}$ is a function of $(\Sigma^{-1})_{11}$. As a result, we expect the $c_{1}$ of the Gaussian BRW with S.I. and N-S.I. jumps to be identical, as confirmed by our numerical results in Figures 7 and 7. We use $p_{0}=0$ for the comparison between the numerical and theoretical results in this section.
@@ 5 Concluding remarks
In this paper, we study the first passage times of an $\mathbb{R}^{d}$-valued branching random walk (BRW) to a shifted unit ball. The first passage time asymptotics consist of a linear term and a logarithmic correction term, as a function of the Euclidean distance of the target ball from the origin. We discuss extensions of this result to the delayed branching BRW model. As an immediate application, we obtain theoretical predictions of the shortest-path statistics for polymer networks consisting of long chains and random cross-links. We conclude with a conjecture on finer asymptotics of the first passage times.
The first passage times of the branching Brownian motion, the continuous-time sibling of the BRW, possess better tractability due to its connection to the Fisher-KPP equation. It is proven in Section 4.1 of [zhang2024modeling] the finer asymptotic for the first passage times of a standard branching Brownian motion that
@@
where the $O_{\mathbb{P}}(1)$ is tight. We refer the readers to [zhang2024modeling] for the proof and a detailed discussion. In view also of the tightness result of Theorem 2, it is natural to postulate the following conjecture for BRW. Recall the constant $\widehat{c}_{1}$ from Section 1.2.
@@
Remark 4 (Status of Conjecture 1).
After the first version of this paper was posted, Conjecture 1 was proved in [blanchet2024tightness]. We retain the conjecture here for historical context and refer the readers to Theorems 1 and 2 of [blanchet2024tightness]. Nevertheless, a stronger form of Conjecture 1 — that the $O_{\mathbb{P}}(1)$ term converges in law — remains unsolved.
{acks}
We thank Amir Dembo, Haotian Gu, and Alexandra Stavrianidi for helpful discussions. We also thank two referees for their detailed comments that significantly improved the presentation of this paper.
{funding}
The material in this paper is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397. Additional support is gratefully acknowledged from NSF 1915967, 2118199, 2229012, 2312204.
@@
Large deviation estimates
@@
Consider $d\geqslant 2$ and a compact set $K\subseteq\mathbb{R}^{d}$. Suppose that the radial component $R$ of the spherically symmetric random variable $\bm{\xi}$} has exponential moments and $\mathbb{P}(R=0)<1$. Let $\bm{\lambda}\in\mathbb{R}^{d}$ be such that $\mathbb{E}_{\mathbb{P}}[e^{\bm{\lambda}\cdot\mathbf{X}}]$ exists and define a probability measure $\mathbb{Q}$ by $\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}(\mathbf{x}):=e^{\bm{\lambda}\cdot\mathbf{x}}/\mathbb{E}_{\mathbb{P}}[e^{\bm{\lambda}\cdot\mathbf{X}}]$ where $\mathbb{P}$ is the law of $\mathbf{X}$. Then the law of $\bm{\xi}$} under $\mathbb{Q}$ satisfies the strong non-lattice property
@@
Moreover, consider an i.i.d. sequence $\{\mathbf{X}_{n}\}_{n\geqslant 1}$ with the same distribution as $\bm{\xi}$} under $\mathbb{Q}$ with a positive definite covariance matrix $V$ and suppose that $\mathbb{E}_{\mathbb{Q}}[\left\lVert{\bm{\xi}}\right\rVert^{d+2}]<\infty$. Then
$\sup_{B}\bigg|\mathbb{P}\Big(\frac{1}{\sqrt{n}}\sum_{j=1}^{n}(\mathbf{X}_{j}-\mathbb{E}_{\mathbb{Q}}[\mathbf{X}_{j}])\in B\Big)-h(B)\bigg|=o(n^{-d/2}),$ (64)
where the supremum is taken over all boxes $B\subseteq K$, $h(B)=(1+o(1))\mu_{\mathbf{0},V}(B)$ as $n\to\infty$, and $\mu_{\mathbf{0},V}$ is the Gaussian measure on $\mathbb{R}^{d}$ with mean $\mathbf{0}$ and covariance $V$.
@@
By Theorem 2.10 of [fang2018symmetric], for $y\in[-1,1]$ and some $C(d)\geqslant 0$,
@@
$z_{R_{1},R_{2}}(r,\bm{\theta})=C(d)r\left(1-\left(\frac{r^{2}-R_{1}^{2}-R_{2}^{2}}{2R_{1}R_{2}}\right)^{2}\right)^{\frac{d-3}{2}},\penalty 10000\ r\in[|R_{1}-R_{2}|,R_{1}+R_{2}],\penalty 10000\ \bm{\theta}\in\mathbb{S}^{d-1}.$
Therefore, the conditional law of $\bm{\zeta}$} under $\mathbb{P}$, and hence also the unconditional law under $\mathbb{Q}$, has a density in $\mathbb{R}^{d}$. If $\mathbb{P}(R=0)>0$, the laws have densities except for an atom at zero, and the densities are non-trivial since $\mathbb{P}(R=0)<1$. By Lemma 4 in Section XV.4 of [feller1991introduction], (63) holds with $\bm{\zeta}$} in place of $\bm{\xi}$}, and hence (63) holds since $\bm{\zeta}$} is a sum of two i.i.d. copies of $\bm{\xi}$}. The final claim follows from Theorem 2(b) of [bhattacharya1978validity], by noting that the sum in (1.14) therein is $o(1)$ as $n\to\infty$ uniformly on compact sets. ∎
@@
Let $\Lambda(\lambda)=\log\phi_{\xi^{(1)}}(\lambda)=\log\mathbb{E}[e^{\lambda{\xi^{(1)}}}]$. It is clear from (A2) that $\xi^{(1)}$} is non-lattice. The Bahadur–Rao theorem (Theorem 3.7.4 of [dembo2009large]) then implies that uniformly in $c\in[-C\log n,C\log n]$,
$\mathbb{P}(F_{n}\geqslant m_{n}+c)\asymp\frac{e^{-nI((m_{n}+c)/n)}}{\sqrt{n}}\asymp\frac{e^{-nI(c_{1})-(c-3\log n/2)I^{\prime}(c_{1})}}{\sqrt{n}}=n\rho^{-n}e^{-c_{2}c}.$ (65)
@@
satisfies that under $\mathbb{Q}$, $\{X_{i}\}_{i\in\mathbb{N}}$ are i.i.d. with mean $c_{1}$. The random variable $(\widetilde{X}_{i},\mathbf{Y}_{i})$ is then centered under $\mathbb{Q}$ where $\widetilde{X}_{i}=X_{i}-c_{1}$. Define $\widetilde{F}_{n}:=\widetilde{X}_{1}+\dots+\widetilde{X}_{n}$. It follows that
$\hskip 14.22636pt\mathbb{P}(F_{n}\geqslant m_{n}+c,\,\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n)))$
$=e^{n\Lambda(c_{2})}\mathbb{E}_{\mathbb{Q}}\left[e^{-c_{2}F_{n}}{\mathbbm{1}}_{\{F_{n}\geqslant m_{n}+c,\,\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n))\}}\right]$
$\asymp e^{n\Lambda(c_{2})-m_{n}c_{2}}\sum_{j=c+1}^{\infty}e^{-c_{2}j}\mathbb{Q}\left(\widetilde{F}_{n}\in\Big[j-1-\frac{3}{2c_{2}}\,\log n,j-\frac{3}{2c_{2}}\,\log n\Big),\,\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n))\right)$
$=\rho^{-n}n^{3/2}\sum_{j=c+1}^{\infty}e^{-c_{2}j}\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in\Big[\frac{j-1-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}},\frac{j-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}}\Big)\times B_{\mathbf{0}}\Big(\frac{u(n)}{\sqrt{n}}\Big)\right).$
@@
$\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in E_{j,n}\right)=(1+o(1))\mu_{\mathbf{0},V}\left(E_{j,n}\right)\asymp u(n)^{d-1}n^{-d/2},$ (67)
@@
$\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n))\mid F_{n}\geqslant m_{n})$ $\ll(n\rho^{-n}e^{-c_{2}c})^{-1}\rho^{-n}n^{3/2}\sum_{j=c+1}^{\infty}e^{-c_{2}j}u(n)^{d-1}n^{-d/2}$
@@
$\hskip 14.22636pt\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{y}}(1),\penalty 10000\ F_{n}\in[m_{n}+c,m_{n}+c+a(n)])$
$\asymp\rho^{-n}n^{3/2}\sum_{j=c+1}^{c+a(n)}e^{-c_{2}j}\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in\Big[\frac{j-1-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}},\frac{j-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}}\Big)\times B_{\frac{\mathbf{y}}{\sqrt{n}}}\Big(\frac{1}{\sqrt{n}}\Big)\right)$
@@
This proves (21). [^14] We have implicitly used that $a(n)\geqslant 1$, but the general case $a(n)\gg 1$ proceeds with the same proof by decomposing $m_{n}+c,m_{n}+c+a(n)]$ into sub-intervals of smaller length than one. The same computation also yields
$\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{0}}(K\sqrt{n}),\penalty 10000\ F_{n}\geqslant m_{n}+c)$ $\ll\rho^{-n}n^{3/2}n^{-1/2}=n\rho^{-n},$